OpenStudy (anonymous):

Town A is 9 miles from town B, and 12 miles from town C. The three towns are NOT collinear (you can’t draw a single straight line through all 3 towns) A road connects towns B and C directly. Find the length of this road.

OpenStudy (anonymous):

there is not enough information in this to know the distance from B to C

OpenStudy (anonymous):

essentially they are talking about a triangle in which we know 2 sides, in order to know the third side, we need an angle

OpenStudy (anonymous):

well that is all that is on the question

OpenStudy (anonymous):

is there a picture?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

well then the answer is anywhere between 3 and 21 miles (not including 3 and 21 as those would be collinear lengths)

OpenStudy (anonymous):

if the AC and AB roads connected at a right angle we would at least be able to use the Pythagorean theorem...

OpenStudy (anonymous):

what have you been working on in class? that could tell us what you are supposed to use

OpenStudy (anonymous):

oh sorry umm the pythagrean thyrum

OpenStudy (anonymous):

oh, then that easy, lets call teh length of the BC road X 9^2+12^2=X^2 81+144=X^2 225=X^2 take the square root sqrt(225)=X 15=X

OpenStudy (anonymous):

thnks, i had to post this 3 times this hour to get THIS ONE answer

OpenStudy (anonymous):

a lot of people wont answer if they think the problem is impossible, which this is without the right angle stuff

OpenStudy (anonymous):

or evven for som1 2 say. their isnt enough info

OpenStudy (anonymous):

well thanks anyway

OpenStudy (anonymous):

glad to help ^^